# TemplateRL: Structured Template-Guided Reinforcement Learning for LLM Reasoning

### v4

@@ Abstract
Reinforcement learning (RL) has emerged as an effective paradigm for enhancing model reasoning. However, existing RL methods like GRPO often rely on unstructured self-sampling to fit scalar rewards, often producing inefficient rollouts that fail to capture transferable problem-solving strategies. To address these limitations, we propose TemplateRL, a structured template-guided RL framework that augments policy optimization with explicit template guidance. Our approach first constructs a problem-solving template library via MCTS on a small seed set, then seamlessly integrates this high-level structured guidance into RL training. By guiding rollout generation to align with proven template structures, TemplateRL significantly improves high-quality trajectory hit rates while reducing ineffective exploration. This structure-guided design steers the policy toward validated strategic patterns, stabilizing training dynamics, and enhancing RL sampling efficiency. Notably, the explicit template library is interpretable, editable, and supports online updates-enabling continuous updates during both training and inference. Extensive experiments demonstrate that TemplateRL outperforms GRPO by 99% on AIME and 41% on AMC, with superior stability on weak models and remarkable cross-domain generalization, highlighting its potential for broader tasks.
@@
Jinyang Wu1thanks:   Equal Contribution, Chonghua Liao [^1] footnotemark: 1, Mingkuan Feng [^1] footnotemark: 1, Shuai Zhang1, Zhengqi Wen5, Haoran Luo2, Ling Yang3, Huazhe Xu1,4, Jianhua Tao1,5 1Tsinghua University, 2Nanyang Technological University, 3Princeton University, 4Shanghai AI Lab, 5Beijing National Research Center for Information Science and Technology {wu-jy23,lch22}@mails.tsinghua.edu.cn haoran.luo@ieee.org, yangling0818@163.com
@@ 3.1 Template Construction
MCTS Exploration.
Starting with a small seed set $\mathcal{S}=\left\{\mathbf{s}_{1},\dots,\mathbf{s}_{s}\right\}$, we employ Monte Carlo Tree Search (MCTS) (Kocsis and Szepesvári, 2006a) to generate solution trees. For each question $\mathbf{s}_{i}\in\mathcal{S}$, given a predefined action set $\mathcal{A}=\{a_{1},\dots,a_{|\mathcal{A}|}\}$ and model $\pi_{\theta}$, MCTS build a search tree $\mathcal{T}_{i}$ where: the root node represents question $\mathbf{s}_{i}$, each edge denotes an action $a\in\mathcal{A}$, and each child node $n$ contains partial solutions generated by $\pi_{\theta}$ under the corresponding action. A path from root $\mathbf{s}_{i}$ to leaf node $\mathbf{n}_{i,j,d_{i,j}}$ forms a solution trajectory:
@@ 4.1 Experimental Setup
For fair comparison, we benchmark TemplateRL with multiple representative baselines, including Standard GRPO (Shao et al., 2024), SimpleRL-Zero (Zeng et al., 2025), OpenReasoner-Zero (Hu et al., 2025b), PRIME-Zero (Cui et al., 2025a), and Oat-Zero (Liu et al., 2025). More details are provided in Appendix D.
@@
Implementation Details.
Following previous work Cui et al. (2025a); Liu et al. (2025), we use Qwen2.5-Math-7B-Base Yang et al. (2024a) as the default model. For RL training, we set $\beta=0$ (no KL loss) and employ Dr.GRPO loss Liu et al. (2025). Our configuration includes batch size 128, 16 samples per question, and 2 guidance templates ($|g|=2$). The reward is binary accuracy verified by Math-Verify. We train for 500 steps on 8 A100 GPUs. More details are provided in Appendix D.
4.2 Improved Performance
Main Results.
@@ 4.2 Improved Performance
As shown in Figure 4, TemplateRL maintains consistent improvements across: (1) different scales, from 1.5B (Qwen2.5-Math-1.5B) to 8B parameters (Llama-3.1-8B-Instruct); (2) diverse architectures, both Qwen and Llama families show substantial improvements; and (3) various model types, both base models and instruction-tuned models benefit from structure-guided training (details in Table 6).
@@ 4.4 Broad Generalization
Beyond textual reasoning, we extend TemplateRL to multimodal domains. We use the Geometry3K dataset (Lu et al., 2021) as training data and evaluate on diverse benchmarks, including mathematical reasoning (MathVision, MathVerse, MathVista), general understanding (MMMU), and visual perception (BLINK). Our settings are: batch size 128, 16 samples per question, and 2 template guidance. As shown in Table 2, TemplateRL consistently outperforms GRPO across all benchmarks, achieving +2.5% average improvement. This indicates that structured template guidance effectively generalizes beyond text modality to multimodal scenarios. More implementation details are provided in Appendix E.3.
@@ 4.5 Dynamic Extensibility
$\bigtriangleup$ $(\uparrow)$ +10.2% +0% +6.6% +4.1%
Training-time Template Expansion.
We explore dynamic library expansion during training by continuously incorporating new templates from successful rollouts. After training on 1k, 2k, and 5k samples, we extract high-level action sequences from correct samples and add them to the library. As shown in Table 3, dynamic expansion achieves substantial improvements across all benchmarks, with notable gains on AIME24 (36.7/33.3=+10.2%). This indicates that our explicit structured library enables continuous updates and reuse of successful exploration experiences, benefiting experience-intensive domains such as medical reasoning where iterative knowledge updates are necessary.
@@
Test-time Template Expansion.
Beyond training, we explore library expansion during inference. For each test sample, we retrieve five templates to generate multiple paths, apply majority voting for the answer, and immediately abstract patterns from voted outputs into the library before processing the next sample. Table 4 shows substantial improvements, demonstrating progressive library enrichment during inference using majority-voted pseudo-supervision. This continuous expansion capability helps for test-time scaling scenarios where models leverage accumulated knowledge from earlier predictions to improve subsequent ones. Note that, while this may introduce some noise, prior work has validated the majority voting’s effectiveness as pseudo-supervision (Qi et al., 2024; Zuo et al., 2025).
@@ 4.6 Ablation Study and Analysis
We provide additional results and analysis in Appendix, including comprehensive OOD evaluation (E.1), detailed extension to multimodal domains and other RL algorithms (E.3,E.4), sensitivity to training data and pattern-matching metrics (E.6,E.7), ablation study on action space (E.9), statistical analysis (E.10), and case study (F).
@@ 5 Related Work
Reinforcement learning (RL) has become a popular paradigm for enhancing LLM reasoning (Guo et al., 2025; Team et al., 2025; Yang et al., 2025). Recent research mainly focuses on: (1) algorithmic refinements that address RL’s inherent limitations (Yu et al., 2025; Liu et al., 2025); (2) data-level optimizations (Zuo et al., 2025; Wang et al., 2025b); and (3) exploration enhancement (Wang et al., 2025a; Cui et al., 2025b). However, these methods remain constrained by unstructured self-sampling, lacking mechanisms to leverage structured guidance (Yue et al., 2025). TemplateRL addresses this by integrating explicit reasoning templates into policy optimization, steering training toward validated strategic patterns.
@@ 6 Conclusion
We introduce TemplateRL, a template-guided RL framework for LLM reasoning. By constructing structured templates and integrating them during policy optimization, TemplateRL steers training toward validated strategic patterns. Extensive experiments demonstrate significant improvements across diverse models, datasets, and modalities. The framework exhibits strong cross-domain generalization, dynamic extensibility, and produces interpretable decision chains, advancing toward more effective and transparent reasoning systems.
@@ C.1 Monte Carlo Tree Search (MCTS)
(1) Selection. Starting from the root node, we identify optimal nodes for expansion by traversing the tree level-by-level until reaching a leaf node (maximum depth or final answer). To balance exploration and exploitation, we employ Upper Confidence Bounds applied to Trees (UCT) (Kocsis and Szepesvári, 2006b):
@@ C.2 Thought Template Library Construction
After MCTS exploration, we obtain multiple solution paths for each question $\mathbf{s}_{i}\in\mathcal{S}$. Following Section 3.1, we identify the best reasoning path and abstract it into generalizable templates. To select the optimal trajectory, we employ the balanced scoring metric:
@@
where $C(\mathbf{t}_{j})$ represents trajectory complexity (action count), and $b=0.95$ balances solution quality against complexity. This function selects trajectories that maximize accuracy while maintaining conciseness.
For each question $\mathbf{s}_{i}\in\mathcal{S}$, we select the optimal trace $\mathbf{t}_{i,\text{best}}=\arg\max_{\mathbf{t}_{j}}\text{Score}(\mathbf{s}_{i},\mathbf{t}_{j})$. Since each node in $\mathbf{t}_{i,\text{best}}$ corresponds to an action $a_{j,\ell}\in\mathcal{A}$, we extract the action sequence as a high-level template $T_{j}=(a_{1},\dots,a_{d})$. For instance, a successful solution trajectory might follow the pattern $a_{1}\rightarrow a_{2}\rightarrow a_{4}$ (System Analysis → One-Step Thought → Divide and Conquer).
To organize extracted patterns, we use Problem Condition Complexity (PCC) (Lee and Heyworth, 2000; Embretson and Daniel, 2008) as a categorization metric. PCC quantifies the number of prior conditions in question $\mathbf{s}_{i}$ and can be computed by model $\pi_{\theta}$. Similar problems tend to share similar PCC values, making this metric effective for pattern aggregation.
Through this process, each question $\mathbf{s}_{i}\in\mathcal{S}$ is associated with its optimal template, with some questions naturally sharing identical patterns. Our final template library $\mathcal{L}=\{\hat{T}_{1},\dots,\hat{T}_{s}\}$ consists of entries where each template $\hat{T}_{j}$ contains both a high-level pattern and the average PCC of questions sharing this pattern: $\hat{T}_{j}=(\text{PCC}_{T_{j}},T_{j})$. These templates represent generalized problem-solving strategies and serve as structured guidance during RL training.
For better visualization, we provide a detailed example thought template in Figure 10.
@@ C.3 Adaptive Template Retrieval
This distance quantifies similarity between the current question and problems from which each template was derived.
@@
We rank templates by distance and select the top-$k$ most similar: $\{\hat{T}_{i_{1}},\dots,\hat{T}_{i_{k}}\}$. These selected templates contain high-level patterns $(a_{1},\dots,a_{d})$ proven effective for problems with similar complexity profiles.
During rollout generation, retrieved templates serve as structured guidance that balances exploiting known successful strategies with model-internal exploration. This adaptive mechanism ensures the model leverages appropriate reasoning strategies based on problem characteristics, enabling more targeted and effective sampling across diverse problem types.

### v5

@@ Abstract
Reinforcement learning (RL) has emerged as an effective paradigm for enhancing model reasoning. However, existing RL methods like GRPO typically rely on unstructured self-sampling to fit scalar rewards, often producing inefficient rollouts that fail to capture transferable problem-solving strategies. To address this limitation, we propose TemplateRL, a structured template-guided RL framework that augments policy optimization with explicit template guidance. Our approach first constructs a problem-solving template library via MCTS on a small seed set, then seamlessly integrates this high-level structured guidance into RL training. By guiding rollout generation to align with proven template structures, TemplateRL significantly improves high-quality trajectory hit rates while reducing ineffective exploration. This structure-guided design steers the policy toward validated strategic patterns, stabilizing training dynamics, and enhancing RL sampling efficiency. Notably, the explicit template library is interpretable, editable, and supports online updates-enabling continuous updates during both training and inference. Extensive experiments demonstrate that TemplateRL outperforms GRPO by 99% on AIME and 41% on AMC, with superior stability on weak models and remarkable cross-domain generalization, highlighting its potential for broader tasks.
@@
Jinyang Wu1thanks:   Equal Contribution, Chonghua Liao [^1] footnotemark: 1, Mingkuan Feng [^1] footnotemark: 1, Shuai Zhang1, Zhengqi Wen5, Haoran Luo2, Ling Yang3, Huazhe Xu1,4, Jianhua Tao1,5 1Tsinghua University, 2Nanyang Technological University, 3Princeton University, 4Shanghai AI Lab, 5Beijing National Research Center for Information Science and Technology {wu-jy23,lch22}@mails.tsinghua.edu.cn
@@ 3.1 Template Construction
1 MCTS Exploration.
Starting with a small seed set $\mathcal{S}=\left\{\mathbf{s}_{1},\dots,\mathbf{s}_{s}\right\}$, we employ Monte Carlo Tree Search (MCTS) (Kocsis and Szepesvári, 2006b) to generate solution trees. For each question $\mathbf{s}_{i}\in\mathcal{S}$, given a predefined action set $\mathcal{A}=\{a_{1},\dots,a_{|\mathcal{A}|}\}$ and model $\pi_{\theta}$, MCTS build a search tree $\mathcal{T}_{i}$ where: the root node represents question $\mathbf{s}_{i}$, each edge denotes an action $a\in\mathcal{A}$, and each child node $n$ contains partial solutions generated by $\pi_{\theta}$ under the corresponding action. A path from root $\mathbf{s}_{i}$ to leaf node $\mathbf{n}_{i,j,d_{i,j}}$ forms a solution trajectory:
@@ 4.1 Experimental Setup
We benchmark TemplateRL with representative baselines, such as Standard GRPO (Shao et al., 2024) and Oat-Zero (Liu et al., 2025). More details are provided in Appendix D.
Implementation Details.
Following previous work Cui et al. (2025a); Liu et al. (2025), we use Qwen2.5-Math-7B-Base Yang et al. (2024a) as the default model. For RL training, we set $\beta=0$ (no KL loss) and employ Dr.GRPO loss Liu et al. (2025). Our configuration includes batch size 128, 16 samples per question, and 2 guidance templates ($|g|=2$). The reward is binary accuracy verified by Math-Verify. We train for 500 steps on 8 A100 GPUs. More details are provided in Appendix D. h
4.2 Improved Performance
Main Results.
@@ 4.2 Improved Performance
As shown in Figure 4, TemplateRL maintains consistent improvements across: (1) different scales, from 1.5B to 8B parameters; (2) diverse architectures, both Qwen and Llama families show substantial improvements; and (3) various model types, both base models and instruction-tuned models benefit from structure-guided training (details in Table 6). We further validate template guidance on larger models (32B) in Appendix E.5, demonstrating that structured guidance enhances rather than restricts exploration even for highly capable models.
@@ 4.4 Broad Generalization
Beyond textual reasoning, we extend TemplateRL to multimodal domains. We use the Geometry3K dataset (Lu et al., 2021) to construct MMTrain10K as training data and evaluate on diverse benchmarks, including mathematical reasoning (MathVision, MathVerse, MathVista), general understanding (MMMU), and visual perception (BLINK). Our settings are: batch size 128, 16 samples per question, and 2 template guidance. As shown in Table 2, TemplateRL consistently outperforms GRPO across all benchmarks, achieving +8.4% average improvement. This indicates that structured template guidance effectively generalizes beyond text modality to multimodal scenarios. More details are provided in Appendix E.3.
@@ 4.5 Dynamic Extensibility
Training-time Template Expansion.
We explore dynamic library expansion during training by continuously incorporating new templates from successful rollouts. After training on 1k, 2k, 3k, and 5k samples, we extract high-level action sequences from correct samples and add them to the library. As shown in Table 3, dynamic expansion achieves substantial improvements across all benchmarks, with notable gains on AIME24 (+10.2%). This indicates that our explicit structured library enables continuous updates and reuse of successful exploration experiences, benefiting experience-intensive domains such as medical reasoning where iterative knowledge updates are necessary.
Test-time Template Expansion.
Beyond training, we explore library expansion during inference. For each test sample, we retrieve five templates to generate multiple paths, apply majority voting for the answer, and immediately abstract patterns from voted outputs into the library before processing the next sample. Table 4 shows substantial improvements, demonstrating progressive library enrichment during inference using majority-voted pseudo-supervision. This continuous expansion capability helps for test-time scaling scenarios where models leverage accumulated knowledge from earlier predictions to improve subsequent ones.
@@
[…]
@@ 4.6 Ablation Study and Analysis
We provide additional results and analysis in Appendix, including comprehensive OOD results (E.1), discussion on template construction strategy (E.6), ablation study (E.12), statistical analysis (E.13), case study (F).
@@ 5 Related Work
Reinforcement learning (RL) has become a popular paradigm for enhancing LLM reasoning (Guo et al., 2025; Team et al., 2025; Tang et al., 2025). Recent research mainly focuses on: (1) algorithmic refinements that address RL’s inherent limitations (Yu et al., 2025; Liu et al., 2025); (2) data-level optimizations (Zuo et al., 2025; Wang et al., 2025b); and (3) exploration enhancement (Wang et al., 2025a; Cui et al., 2025b). However, these methods remain constrained by unstructured self-sampling, lacking mechanisms to leverage structured guidance (Yue et al., 2025). TemplateRL addresses this by integrating explicit reasoning templates into policy optimization, steering training toward validated strategic patterns.
@@ 6 Conclusion
We introduce TemplateRL, a template-guided RL framework for LLM reasoning. By constructing structured templates and integrating them during policy optimization, TemplateRL steers training toward validated strategic patterns. Experiments demonstrate significant improvements across models, datasets, and modalities. The framework exhibits cross-domain generalization, dynamic extensibility, and produces interpretable decision chains.
@@
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. U2436210, No. 62322120).
@@ C.1 Monte Carlo Tree Search (MCTS)
(1) Selection. Starting from the root node, we identify optimal nodes for expansion by traversing the tree level-by-level until reaching a leaf node (maximum depth or final answer). To balance exploration and exploitation, we employ Upper Confidence Bounds applied to Trees (UCT) (Kocsis and Szepesvári, 2006a):
@@ C.2 Thought Template Library Construction
After MCTS exploration, we obtain multiple solution paths for each question $\mathbf{s}_{i}\in\mathcal{S}$. We then construct a thought template library, which can be used in RL optimization and update during both training and testing process.
Template Construction Process.
The construction involves five key steps:
  1. 1.
    Seed Set Selection: We select a small set of seed problems $\mathcal{S}=\{s_{1},s_{2},\dots,s_{s}\}$ (500 samples in our experiments). These problems are chosen to be complex, while also providing diverse coverage of possible solution strategies.
  2. 2.
    MCTS Exploration: For each seed problem $s_{i}\in\mathcal{S}$, we apply MCTS to explore the solution space. MCTS recursively simulates different action sequences and evaluates resulting solution paths, generating a solution tree $\mathcal{T}_{i}$ where each path represents a candidate reasoning trajectory.
  3. 3.
    Optimal Path Selection: From the multiple solution paths generated by MCTS, we select the best trajectory for each question using a balanced scoring metric:
@@
    where $R(\mathbf{t}_{j}|\mathbf{s}_{i})$ is the reward, $C(\mathbf{t}_{j})$ is trajectory complexity (action count), and $b=0.95$ balances solution quality against conciseness. The optimal trace, $\mathbf{t}_{i,\text{best}}$, is selected as the one that maximizes this score:
    $\mathbf{t}_{i,\text{best}}=\arg\max_{\mathbf{t}_{j}}\text{Score}(\mathbf{s}_{i},\mathbf{t}_{j}).$
  4. 4.
    Template Abstraction: Since each node in $\mathbf{t}_{i,\text{best}}$ corresponds to an action $a_{\ell}\in\mathcal{A}$, we extract the action sequence as a high-level template $T_{i}=(a_{1},\dots,a_{d})$. For instance, a successful solution might follow the pattern $a_{1}\rightarrow a_{2}\rightarrow a_{4}$ (System Analysis → One-Step Thought → Divide and Conquer). Thus, a template is defined as an abstract, high-level sequence of actions that represents a strategic problem-solving approach to a particular task.
  5. 5.
    Template Aggregation: To organize the extracted patterns, we use Problem Condition Complexity (PCC) (Lee and Heyworth, 2000; Embretson and Daniel, 2008) as a categorization metric. PCC quantifies the structural complexity of a problem based on the number of prior conditions and constraints, which can be computed by prompting model $\pi_{\theta}$ with the question. Templates sharing similar PCC values are grouped together, as problems with comparable structural complexity tend to benefit from similar reasoning strategies. Through this aggregation, each template $\hat{T}_{j}$ stores both its action pattern and the average PCC of questions sharing this pattern:
    $\hat{T}_{j}=(\text{PCC}_{T_{j}},T_{j}),$
    where $\text{PCC}_{T_{j}}=\frac{1}{|I_{j}|}\sum_{i\in I_{j}}\text{PCC}(s_{i})$ and $I_{j}=\{i:T_{i}=T_{j}\}$.
Final Template Library.
The final template library $\mathcal{L}=\{\hat{T}_{1},\dots,\hat{T}_{m}\}$ consists of 103 unique templates derived from 500 seed samples. Each template represents a generalized problem-solving strategy characterized by both its structural pattern (action sequence) and typical problem complexity (average PCC). These templates serve as structured guidance during RL training, enabling the policy to leverage proven strategic patterns when encountering similar problems. The final library thus provides a powerful tool for both training and dynamic adaptation during inference.
For better visualization, we provide a detailed example template in Figure 10.
Figure 11: Concrete example of adaptive template retrieval. Given a 3D distance calculation problem with $\text{PCC}=2$, the system retrieves top-2 templates ($T_{1}$ and $T_{2}$) based on complexity similarity while filtering out the overly complex $T_{3}$ ($d=1.5$). The retrieved templates guide distinct reasoning trajectories: $T_{1}$ provides a streamlined solution path, while $T_{2}$ incorporates an additional verification step for enhanced reliability.
@@ C.3 Adaptive Template Retrieval
This distance quantifies similarity between the current question and problems from which each template was derived. Note that, the PCC prompt is:
You are an AI assistant to help me rephrase questions by splitting the question context into conditions. In your rephrased question, remember to fully express the information in the original question. That means, you should count the number of distinct prior conditions or given facts that must be satisfied or considered to solve it.
Problem: <problem>
Number of prior conditions:
@@
We rank templates by distance and select the top-$k$ most similar: $\{\hat{T}_{i_{1}},\dots,\hat{T}_{i_{k}}\}$. These selected templates contain high-level patterns $(a_{1},\dots,a_{d})$ proven effective for problems with similar complexity profiles. During rollout generation, retrieved templates serve as structured guidance that balances exploiting known successful strategies with model-internal exploration. This adaptive mechanism ensures the model leverages appropriate reasoning strategies based on problem characteristics, enabling more targeted and effective sampling across diverse problem types.
Concrete Example.
We provide a detailed processing example in Figure 11. This example demonstrates how PCC-based retrieval identifies templates with appropriate complexity levels (rejecting the overly complex $T_{3}$), and how different templates induce varied reasoning structures. Template $T_{1}$ provides a concise solution path, while $T_{2}$ incorporates an additional verification step that enhances answer reliability—both are valid strategies, but $T_{2}$ may be preferred in high-stakes scenarios where correctness verification is critical. During RL training, the model learns to leverage these structural variations to maximize expected rewards across diverse problem instances.
Figure 12: The system prompt used for all experiments.
Template Updates.
Regarding template updates, during RL training, we optionally leverage correct rollouts to continually enrich the template library. For each correct rollout, we apply keyword-based or lightweight model-based pattern extraction to automatically identify the underlying action sequence (template), which is then incorporated into the library for continuous expansion. Tables 3-4 empirically validate the effectiveness of dynamic updates during both training and inference.
C.4 Discussion on Additional Method Details
This section further clarifies several key details regarding template construction and abstraction.
Template Aggregation via Exact Action Sequence Matching.
Templates are aggregated through exact matching. Two solution traces share the same template if and only if they follow identical action sequences. For example, all traces following $a_{1}\rightarrow a_{4}\rightarrow a_{3}$” are merged into a single template. The aggregation process treats each unique sequence $T$ as a key, collects all questions matching $T$ with their PCC values, and computes the average PCC to obtain the template’s characteristic complexity score, storing the $(T,\text{PCC})$ pair. This exact matching ensures templates represent genuinely identical reasoning structures.
Predefined Actions vs. Automatically Discovered Patterns.
Our framework employs a two-level design: The action set $\mathcal{A}$ (e.g., divide-and-conquer) comprises predefined atomic reasoning primitives. However, templates (sequences of these actions) are automatically discovered through MCTS exploration. For instance, while $a_{1}$” (divide-and-conquer) is predefined, the template $a_{1}\rightarrow a_{4}\rightarrow a_{3}$” emerges from successful solution paths. This design provides interpretability through recognizable reasoning concepts while enabling discovery of effective strategy combinations that may not be obvious to human experts.
@@ C.4 Discussion on Additional Method Details
Robustness to Paraphrasing through Action-Level Abstraction.
Template matching operates at the action level rather than natural language level. During MCTS, each reasoning step is generated by a specific action prompt and assigned the corresponding action label. Different surface expressions of the same reasoning step (e.g., “break into x-y coordinates” vs. “decompose dimension-wise” vs. “handle each axis separately”) are unified through their shared action label $a_{1}$. When abstracting traces into templates, we extract action sequences rather than literal text, making templates inherently robust to paraphrasing. This contrasts with approaches that directly cluster natural language trajectories, which often require complex similarity metrics to handle linguistic variations.